OpenStudy (anonymous):

Which expression shows the value of sec x°? A) adjacent/ opposite B) adjacent/ hypotenuse C) hypotenuse/ opposite D) hypotenuse/ adjacent

OpenStudy (anonymous):

secant is opposite of cosine. .

OpenStudy (anonymous):

sec=1/cos

OpenStudy (anonymous):

cos=a/h

OpenStudy (anonymous):

so... is h/a :)

OpenStudy (anonymous):

d.

OpenStudy (lgbasallote):

CHO SHA CAO cosecant = H/O secant = hypotenuse/adjacent cotangent = adjacent/opposite additional info SOHCAHTOA sin = opp/hyp cos = adj/hyp tan = opp/adj

OpenStudy (anonymous):

(: Thankkss